OpenStudy (anonymous):

What is F'(x) when F(x)=Integral from 0 to x of (sin^2(t) dt)

OpenStudy (anonymous):

\[F(x)=\int\limits_{0}^{x}\sin^2(t)dt\] What is F'(x)

OpenStudy (anonymous):

you need to use a trig identity for sin^2(t)

OpenStudy (anonymous):

what is a trig identity?

OpenStudy (anonymous):

you are in calculus but you do not know what a trig identity is?

OpenStudy (anonymous):

im bad with terminology. I thought that the derivative of sin^2(t) was just 2cost

OpenStudy (anonymous):

\[\sin^{2}(x) = \frac{1}{2}(1 + \cos(2t))\]

OpenStudy (anonymous):

you can replace sin with the right side of the equality

OpenStudy (anonymous):

ok but what about the fact that it's F(x) and it's t in the function?

OpenStudy (anonymous):

just replace the x with a t

OpenStudy (anonymous):

oh lol ok

OpenStudy (anonymous):

i didn't know that thx

OpenStudy (anonymous):

\[F(x)=\int\limits\limits_{0}^{x}\frac{1}{2}(1 + \cos(2t))dt\]

OpenStudy (anonymous):

\[F(x)=\frac{1}{2}\int\limits\limits\limits_{0}^{x}(1 + \cos(2t))dt\]

OpenStudy (anonymous):

so \[F'(x)=\int\limits_{0}^{x}2\cos(2t)\] Right?

OpenStudy (anonymous):

no you need to integrate the first function

OpenStudy (anonymous):

so \[F(x)=1/2(t-\sin(2t)/2)\]

OpenStudy (anonymous):

yeah no plug in the upper and lower limits

OpenStudy (anonymous):

yeah i guess my identiy was a little of but it looks like you corrected it

OpenStudy (anonymous):

ok so then now i can find f'(x) right?

OpenStudy (anonymous):

which would be 1/2-cos(2t)/2 right?

OpenStudy (anonymous):

yes but you were supposed to plulg in the upper and lower limits which would of changed it to x

OpenStudy (anonymous):

ok, so it would be\[F'(x)=(1-\cos(2x))\div2\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thx